OpenStudy (gabylovesyou):

The function for the cost of materials to make a muffin is f(x) = Three fourths x + 3, where x is the number of muffins. The function for the selling price of those muffins is g(f(x)), where g(x) = 3x + 4. Find the selling price of 20 muffins. 18 49 51 58

OpenStudy (neonumbrella5115):

I made a mistake since it is g of fx then you will need to put the fx equation into the g equation as the x and then solve for x.

OpenStudy (gabylovesyou):

sooo.... g(3/4 *18 + 3) = 3(3/4*18+3) + 4

OpenStudy (gabylovesyou):

@ganeshie8 @hartnn

OpenStudy (neonumbrella5115):

Correct! I would simplify and then plug in the 20! You're good on simplifying?

OpenStudy (gabylovesyou):

i have no idea how to do this part now... .-.

OpenStudy (solomonzelman):

Your f(x) is, \(\color{#000000 }{ \displaystyle f(x)=\frac{3}{4}(x+3) }\) or is it, \(\color{#000000 }{ \displaystyle f(x)=\frac{3}{4}x+3 }\)

OpenStudy (gabylovesyou):

no parenthesis :) so first.

OpenStudy (gabylovesyou):

i mean second lmao

OpenStudy (solomonzelman):

The selling price any "x" muffins is modeled as \(\color{#000000 }{ \displaystyle g(f(x))=4(f(x))+4=4(\frac{3}{4}x+3)+3 }\) ($ of selling muffins) So, selling price of 20 muffins would be \(\color{#000000 }{ \displaystyle g(f(\color{red}{20}))=3(\frac{3}{4}\cdot \color{red}{20}+3)+4=~.... }\)

OpenStudy (gabylovesyou):

idk how to simplify ;( @SolomonZelman

OpenStudy (gabylovesyou):

is it 58 ?

OpenStudy (solomonzelman):

I would simplify at this point, \(\color{#000000 }{ \displaystyle g(f(x))=4(f(x))+4 =4(\frac{3}{4}x+3)+3\\[1.3em] \displaystyle {\small \rm (here)\quad \quad } =(3x+12)+3=3x+15 }\)

OpenStudy (solomonzelman):

you shoud be able to tell that for even x, you get even+odd=odd, so 58 is incorrect.

OpenStudy (solomonzelman):

oh, sorry

OpenStudy (solomonzelman):

I would simplify at this point, \(\color{#000000 }{ \displaystyle g(f(x))=4(f(x))+4 =4(\frac{3}{4}x+3)+4\\[1.3em] \displaystyle {\small \rm (here)\quad \quad } =(3x+12)+4=3x+16 }\)

OpenStudy (solomonzelman):

but, 58 is still not correct

OpenStudy (solomonzelman):

(although it would actually be even)

OpenStudy (solomonzelman):

3•20+16 =?

OpenStudy (gabylovesyou):

76

OpenStudy (gabylovesyou):

@SolomonZelman